OpenStudy (anonymous):

What is the solution of |2x+3| ≥ 18?

OpenStudy (crashonce):

both sides minus 3 then halve

OpenStudy (anonymous):

equation

OpenStudy (anonymous):

how

OpenStudy (crashonce):

\[2x \ge 15\]

OpenStudy (crashonce):

\[x \ge 7.5\]

OpenStudy (anonymous):

this the answer

OpenStudy (anonymous):

explain

OpenStudy (crashonce):

both sides -3

OpenStudy (crashonce):

then voth divide by 2

OpenStudy (anonymous):

the answer is x≥7.5

OpenStudy (crashonce):

yes

OpenStudy (anonymous):

one question

OpenStudy (anonymous):

(2x+5/x-2)=5

OpenStudy (crashonce):

times both sides by x-2: 2x+5 = 5x-2

OpenStudy (crashonce):

rearrange: 3x = 7 x = 7/3

OpenStudy (anonymous):

wait crash

OpenStudy (anonymous):

one last question

OpenStudy (crashonce):

ya

OpenStudy (anonymous):

|3x-5| ≤ 4?

OpenStudy (crashonce):

add 5: \[3x \le 9\]

OpenStudy (crashonce):

divide by 3

OpenStudy (crashonce):

\[x \le 3\]

OpenStudy (anonymous):

:) thanks